# BP-Seg: A graphical model approach to unsupervised and non-contiguous text segmentation using belief propagation

### v1

@@ Abstract
BP-Seg: A graphical model approach to unsupervised and non-contiguous text segmentation using belief propagation
Fengyi Li, Kayhan Behdin, Natesh Pillai, Xiaofeng Wang, Zhipeng Wang, Ercan Yildiz LinkedIn Corporation {fenli, kbehdin, napillai, xiaofwan, zhipwang, eyildiz}@linkedin.com
@@ 1 Introduction
Segmenting text into semantically coherent segments has been a long-studied problem in the field of natural language processing (Pak and Teh, 2018; Badjatiya et al., 2018). The applications of text segmentation range from information retrieval (Yu et al., 2023; Huang et al., 2003), document summarization (Cho et al., 2022; Miculicich and Han, 2023), disclosure analysis (Wang et al., 2018; Song, 2021), and optimizing prompts for large language models (LLMs) by extracting the most relevant parts (Lewis and Brown, 2023).
@@
Traditional methods, whether supervised (Koshorek et al., 2018; Badjatiya et al., 2018; Glavas and Somasundaran, 2020) or unsupervised (Hearst, 1997; Glavaš et al., 2016; Barakat et al., 2020), mainly focus on contiguous or sequential text segmentation. The goal is to cluster consecutive sentences in a way that ensures those within the same group are semantically more similar to each other than to sentences in different groups. For example, if a text consists of five sentences labeled $\{1,2,3,4,5\}$, traditional methods might segment them into groups such as $\{1,2\}$ and $\{3,4,5\}$. However, in practice, it is sometimes the case that $\{1,5\}$ are more semantically similar and should form one group, while $\{2,4,5\}$ form another. On the other hand, some frameworks for text segmentation disregard the adjacency relationships between sentences and the overall structure of the text. For example, methods such as $k$-means (Lloyd, 1982) treat each sentence (or its embedding) as an isolated data instance, without considering that a sentence is often more likely to be semantically connected to its adjacent sentences than to those that are farther apart. To the best of our knowledge, however, there is a lack of literature on semantic classification methods that take into account both adjacent and distant (non-adjacent) sentences. One example of such an application is prompt pruning for LLMs. When users write prompts, the sentences typically follow a logical flow, but some may be redundant. Splitting a prompt into groups, potentially non-sequential, while still accounting for the semantic coherence of contiguous sentences can facilitate downstream tasks such as prompt pruning, ultimately improving both efficiency and relevance in LLM interactions (Gao et al., 2024).
@@
In this work, we propose a new framework for text segmentation that accounts for the fact that adjacent sentences are typically more related, while also enabling the grouping of non-contiguous sentences that are semantically similar. To achieve this, we first embed sentences into vector representations using sentence embeddings Reimers and Gurevych (2019), so that semantically similar sentences are mapped closer together in the embedding space. This allows us to form a graph from the text, where the nodes represent the embedded sentences and edges encode the strength of their semantic relationships. We then apply Belief Propagation (BP, Pearl (1982)), an inference algorithm used in graphical models, to generate clusters. To the best of our knowledge, this work presents the first successful application of BP to text segmentation that accounts for semantic meaning of sentences in both continuous and non-continuous settings.
@@ 2.2 Constructing the graphical model
$\psi_{i,j}(x_{i},x_{j})=\begin{cases}1,&x_{i}=x_{j}\\ \exp(-\phi_{i,j})&\mathrm{otherwise,}\end{cases}$ (2)
where
$\phi_{i,j}=\frac{\lambda}{2}\left(-\mathrm{sim}(R_{i},R_{j})+1\right)\exp(-|i-j|^{2}/\sigma).$ (3)
Here, $\mathrm{sim}(\cdot,\cdot)$ denotes the cosine similarity between two embeddings. A higher value of $\psi_{i}(x_{i})$ indicates a greater likelihood that sentence $S_{i}$ belongs to segment $x_{i}$. Note also that $\psi_{i,j}(x_{i},x_{j})\leq 1$, with equality holding if and only if $x_{i}=x_{j}$, i.e., when the two sentences are assigned the same label. A larger $\phi_{i,j}$ leads to a smaller $\psi_{i,j}$, indicating that semantically dissimilar sentences possess weaker connections.
Although in this work, we adopt specific forms of $\psi_{i}$ and $\psi_{i,j}$ to encode the semantic relationships between sentences and their assignments. However, in practice, one may choose alternative or domain-specific formulations, provided they are compatible with the desired inference algorithm.
@@ 2.3 BP (Sum-Product) for text segmentation
$m_{i\to j}^{(t)}(x_{j})\;=$ $\;\sum_{x_{i}}\Big{(}\psi_{i}(x_{i})\;\psi_{i,j}(x_{i},x_{j})\Big{.}$
.$\Big{.}\prod_{k\in\{1,\ldots,n\})\setminus j}m_{k\to i}^{(t-1)}(x_{i})\Big{)},$ (4)
@@
We summarize our proposed algorithm in Algorithm 1. More analysis of BP can be read in Murphy et al. (2013); Yedidia et al. (2003).
@@ 3 Related work
A prior work has explored a variation of using graph-based models for text segmentation. GraphSeg, proposed by Glavaš et al. (2016), for example, also employs unsupervised learning for text segmentation within a graph-based framework. However, their primary objective is to produce contiguous segmentations, whereas our method allows for a non-contiguous segmentation that accounts for both neighboring and distant sentences. Moreover, their algorithm requires additional information, such as each word’s information content based on its relative frequency, whereas BP-seg relies solely on embeddings and no external data. Moreover, after encoding sentences into embeddings and computing cosine similarities, their approach discovers segmentations by finding maximal cliques — fundamentally different from our probabilistic strategy, in which we seek an assignment that maximizes the marginal distribution. One could in principle apply $k$-means to group sentence embeddings; however, such a method is entirely context-agnostic and considers only pairwise embedding similarities.
@@
Table 1: Average performance on the Choi dataset measured using ARI and NMI across different subsets. Higher values indicate better performance. Note that a random segmentation method achieves a 00 in ARI.
@@ 4.1 Illustrative Example
Each sentence is treated as an individual text unit for segmentation. We analyze the segmentation results obtained using our proposed approach BP-Seg, GraphSeg, $k$-means, and an LLM (GPT-4o). To evaluate the results, we verify (1) whether the algorithm produces more than one cluster, and (2) whether the three tennis-related sentences colored in red appear together in the same cluster. The detailed experimental details and results with various parameters are presented in Appendix A.1.
For BP-Seg and $k$-means, we set the number of segments to $k={2,3,\cdots,7}$. For the LLM, we explicitly prompt it to generate $k={2,3,\cdots,7}$ segments. The prompt can be read in A.1.4. As observed, all methods generate more than one clusters. Furthermore, $k$-means successfully group sentences based on thematic coherence across all values of $k$. Both BP-Seg and the LLM successfully group the tennis-related sentences in the same cluster in $5$ out of the $6$ tested values of $k$. For GraphSeg, we set the minimum number of sentences per segment to 1 to allow maximum flexibility. However, despite trying various thresholds $\tau$, GraphSeg fails to cluster thematically related sentences effectively, as it is designed for contiguous segmentation.
@@ 4.2 Choi dataset
In this example, we implement our approach, BP-Seg, along with GraphSeg and $k$-means on the Choi dataset (Choi, 2000). The performance of the LLM is not included in this case, as its output does not include every sentence from the original text. Traditionally, for contiguous text segmentation, two evaluation metrics are commonly reported: $P_{k}$ (Beeferman et al., 1999) and WindowDiff (WD) (Pevzner and Hearst, 2002). The $P_{k}$ metric checks whether the boundary status (i.e., whether two sentences within a fixed-size window belong to the same segment) matches between the ground truth and the prediction. WD, on the other hand, measures whether the number of boundaries within the window is consistent with the ground truth. However, both metrics assume contiguity and are not suitable for evaluating non-contiguous text segmentation.
Therefore, we report Adjusted Rand Index (ARI) (Hubert and Arabie, 1985; Wagner and Wagner, 2007) and Normalized Mutual Information (NMI) (Kvålseth, 2017), which are appropriate for clustering-based evaluations. This adjustment accounts for the fact that the outputs of BP-Seg and $k$-means may result in non-contiguous segmentations, even though the ground truth segmentation is contiguous. Additionally, for efficiency, we use a variant of BP-Seg that is also based on message passing but offers faster computation. Please refer to Algorithm 2 for more details.
The Adjusted Rand Index (ARI) ranges from $-1$ to $1$, where $1$ indicates a perfect match between predicted and true clusters, 00 corresponds to random labeling, and negative values indicate performance worse than random. The Normalized Mutual Information (NMI) ranges from 00 to $1$, with $1$ representing perfect alignment and 00 indicating statistical independence. Table 2 presents the average segmentation performance of the three methods on the Choi dataset. As shown, BP-Seg outperforms the other two methods across all subsets, except in cases where each segment contains very few sentences (e.g., 3–5). The standard deviation of the performance can be read in Table 2.
@@ 5 Conclusion
We presented BP-Seg, an efficient unsupervised approach for text segmentation using belief propagation. Our method effectively balances local contextual coherence with global semantic similarity, enabling more meaningful and flexible segmentation of text. Although designed for non-contiguous segmentation, experimental results show that BP-Seg outperforms several competitive methods on the standard contiguous segmentation task, achieving strong performance on metrics such as ARI and NMI. In future work, we aim to evaluate the effectiveness of our approach in real-world downstream applications, including LLM prompt pruning, information retrieval, and question answering.
@@
In this study, all examples are in English. The example in Section 4.1 was generated by GPT-4o with additional human-written content, and the Choi dataset in Section 4.2 is also synthetic. As a result, these examples may not accurately reflect real-world scenarios, and our evaluations are limited to these two cases. Nevertheless, we believe the insights from our findings will inspire further research in text segementation and benefit a wide range of related applications.
Ethics Statement
In this paper, we utilize data from two sources: one generated by ChatGPT with additional human-written sentences, and another from a publicly available dataset. We anticipate that our method can offer benefits such as enhanced information retrieval and reduced token usage in LLM prompts, thereby reaching a broader audience.
@@ A.1.1 Results using BP-Seg
: [Segment 1]: The sun was shining brightly. It was a beautiful morning. I decided to go for a walk. Suddenly, dark clouds appeared. I’ll play tennis tomorrow. What are you doing? People ran for shelter. US Open is a tennis tournament. I am here working on my project. The sun came out again. Who is going to win the US Open? [Segment 2]: Thunder rumbled in the distance. The rain poured down heavily.
@@
$k=4$ segments: [Segment 1]: The sun was shining brightly. It was a beautiful morning. What are you doing? People ran for shelter. I am here working on my project. The sun came out again. [Segment 2]: I decided to go for a walk. Suddenly, dark clouds appeared. The rain poured down heavily. [Segment 3]: Thunder rumbled in the distance. [Segment 4]: I’ll play tennis tomorrow. US Open is a tennis tournament. Who is going to win the US Open?
$k=5$ segments: [Segment 1]: The sun was shining brightly. It was a beautiful morning. Suddenly, dark clouds appeared. People ran for shelter. US Open is a tennis tournament. The sun came out again. [Segment 2]: I decided to go for a walk. [Segment 3]: I’ll play tennis tomorrow. What are you doing? Who is going to win the US Open? [Segment 4]: Thunder rumbled in the distance. The rain poured down heavily. [Segment 5]: I am here working on my project.
$k=6$ segments: [Segment 1]: The sun was shining brightly. It was a beautiful morning. The sun came out again. [Segment 2]: I decided to go for a walk. [Segment 3]: I’ll play tennis tomorrow. US Open is a tennis tournament. Who is going to win the US Open? [Segment 4]: Suddenly, dark clouds appeared. Thunder rumbled in the distance. The rain poured down heavily. [Segment 5]: People ran for shelter. [Segment 6]: What are you doing? I am here working on my project.
@@ A.2 Choi dataset
Here $m_{i}(x_{i})$ can be viewed as a running preference or partial belief for node $i$ in label $x_{i}$. We also set $\psi_{i}(x_{i})=\mathrm{sim}(R_{i},C_{x_{i}})$ and $\psi_{i,j}=\frac{\lambda}{2}\left(-\mathrm{sim}(R_{i},R_{j})+1\right)\exp(-|i-j|^{2}/\sigma)$. For comparison, we set the number of initial clusters in BP-Seg to be equal to the number of sentences, although many clusters may become empty after the algorithm terminates. We fix the number of iterations to $T=5$, and choose $\sigma=10$ and $\lambda=300$ to account for the length of the text. For GraphSeg, we set the threshold parameter to $\tau=0.2$ and the minimum segment size to $n=1$ to allow maximum flexibility. For the $k$-means baseline, we cap the number of clusters $k$ at 20 and rely on the default number of iterations in sklearn.cluster.KMeans.

### v2

# Structured Probabilistic Inference & Generative Modeling

@@ 1 Introduction
Segmenting text into semantically coherent segments has been a long-studied problem in the field of natural language processing [18, 1]. The applications of text segmentation range from information retrieval [27, 10], document summarization [4, 16], disclosure analysis [24, 22], and optimizing prompts for large language models (LLMs) by extracting the most relevant parts [14, 25].
@@
Traditional methods, whether supervised [12, 1, 7] or unsupervised [9, 8, 2], mainly focus on contiguous or sequential text segmentation. The goal is to cluster consecutive sentences in a way that ensures those within the same group are semantically more similar to each other than to sentences in different groups. For example, if a text consists of five sentences labeled $\{1,2,3,4,5\}$, traditional methods might segment them into groups such as $\{1,2\}$ and $\{3,4,5\}$. However, in practice, it is sometimes the case that $\{1,5\}$ are more semantically similar and should form one group, while $\{2,3,4\}$ form another. On the other hand, some frameworks for text segmentation disregard the adjacency relationships between sentences and the overall structure of the text. For example, methods such as $k$-means [15] treat each sentence (or its embedding) as an isolated data instance, without considering that a sentence is often more likely to be semantically connected to its adjacent sentences than to those that are farther apart. To the best of our knowledge, however, there is a lack of literature on semantic classification methods that take into account both adjacent and distant (non-adjacent) sentences. One example of such an application is prompt pruning for LLMs. When users write prompts, the sentences typically follow a logical flow, but some may be redundant. Splitting a prompt into groups, potentially non-sequential, while still accounting for the semantic coherence of contiguous sentences can facilitate downstream tasks such as prompt pruning, ultimately improving both efficiency and relevance in LLM interactions [6, 25].
@@
In this work, we propose a new framework for text segmentation that accounts for the fact that adjacent sentences are typically more related, while also enabling the grouping of non-contiguous sentences that are semantically similar. To achieve this, we first embed sentences into vector representations using sentence embeddings [21], so that semantically similar sentences are mapped closer together in the embedding space. This allows us to form a graph from the text, where the nodes represent the embedded sentences and edges encode the strength of their semantic relationships. We then apply Belief Propagation (BP, Pearl [19]), an inference algorithm used in graphical models, to generate clusters. To the best of our knowledge, this work presents the first successful application of BP to text segmentation that accounts for semantic meaning of sentences in both continuous and non-continuous settings.
@@ 2.2 Constructing the graphical model
$\psi_{i,j}(x_{i},x_{j})=\begin{cases}1,&x_{i}=x_{j}\\ \exp(\lambda\left(\mathrm{sim}(R_{i},R_{j})-1\right))&\mathrm{otherwise}\end{cases}.$ (2)
Here, $\mathrm{sim}(\cdot,\cdot)$ denotes the cosine similarity between two embeddings. A higher value of $\psi_{i}(x_{i})$ indicates a greater likelihood that sentence $S_{i}$ belongs to segment $x_{i}$. Note also that $\psi_{i,j}(x_{i},x_{j})\leq 1$, with equality holding if and only if $x_{i}=x_{j}$, i.e., when the two sentences are assigned the same label. On the other hand, smaller values of $\psi_{i,j}$ reflect weaker semantic connections between sentences that differ in meaning.
In this work, we adopt specific forms of $\psi_{i}$ and $\psi_{i,j}$ to encode the semantic relationships between sentences and their assignments. However, in practice, one may choose alternative or domain-specific formulations, provided they are compatible with the desired inference algorithm.
@@ 2.3 BP (Sum-Product) for text segmentation
$m_{i\to j}^{(t)}(x_{j})=\sum_{x_{i}}\Big(\psi_{i}(x_{i})\;\psi_{i,j}(x_{i},x_{j})\prod_{k\in\{1,\ldots,n\}\setminus j}m_{k\to i}^{(t-1)}(x_{i})\Big),$ (3)
@@
We summarize our proposed algorithm in Algorithm 1. More analysis of BP can be read in Murphy et al. [17], Yedidia et al. [26].
@@ 3 Related work
A prior work has explored a variation of using graph-based models for text segmentation. GraphSeg, proposed by Glavaš et al. [8], for example, also employs unsupervised learning for text segmentation within a graph-based framework. However, their primary objective is to produce contiguous segmentations, whereas our method allows for a non-contiguous segmentation that accounts for both neighboring and distant sentences. Moreover, their algorithm requires additional information, such as each word’s information content based on its relative frequency, whereas BP-seg relies solely on embeddings and no external data. Moreover, after encoding sentences into embeddings and computing cosine similarities, their approach discovers segmentations by finding maximal cliques — fundamentally different from our probabilistic strategy, in which we seek an assignment that maximizes the marginal distribution. One could in principle apply $k$-means to group sentence embeddings; however, such a method is entirely context-agnostic and considers only pairwise embedding similarities.
@@
Table 1: Average performance on the Choi dataset measured using ARI and NMI across different subsets. Higher values indicate better performance. Note that a random segmentation method achieves a 0 in ARI.
@@ 4.1 The Illustrative Example
Each sentence in the document is treated as an individual unit for segmentation. We evaluate the segmentation outcomes generated by our proposed method, BP-Seg, in comparison with three baselines: GraphSeg, $k$-means clustering, and a LLM (GPT-4o). To qualitatively assess the results, we focus on two key criteria: (1) whether the method identifies more than one meaningful segment (i.e., produces more than one cluster), and (2) whether the three semantically related tennis sentences, highlighted in red, are grouped together in the same cluster. These criteria help determine both the model’s ability to detect structure and its sensitivity to semantic coherence. The full experimental setup, including parameter choices and additional visualizations across methods and configurations, can be found in Appendix A.1.
For BP-Seg and $k$-means, we set the number of segments to $k={2,3,\cdots,7}$. For the LLM, we explicitly prompt it to generate $k={2,3,\cdots,7}$ segments. The prompt can be read in A.1.4. As observed, all methods generate more than one clusters. Furthermore, $k$-means successfully group sentences based on thematic coherence across all values of $k$. Both BP-Seg and the LLM successfully group the tennis-related sentences in the same cluster in $5$ out of the $6$ tested values of $k$. For GraphSeg, we set the minimum number of sentences per segment to 1 to allow maximum flexibility. However, despite trying various thresholds $\tau$, GraphSeg fails to cluster semantically related sentences effectively, as it is designed for contiguous segmentation. We summarize the results in Table 2.
2 Clusters 3 Clusters 4 Clusters 5 Clusters 6 Clusters 7 Clusters
BP-Seg
$k$-means
LLM
Graph-Seg
Table 2: Performance on an illustrative example. Both BP-Seg and $k$-means are able to group tennis-related sentences into the same clusters, with varying numbers of clusters.
@@ 4.2 Choi dataset
In this example, we implement our approach, BP-Seg, along with GraphSeg and $k$-means on the Choi dataset [5]. The performance of the LLM is not included in this case, as its output does not include every sentence from the original text. Traditionally, for contiguous text segmentation, two evaluation metrics are commonly reported: $P_{k}$ [3] and WindowDiff (WD) [20]. The $P_{k}$ metric checks whether the boundary status (i.e., whether two sentences within a fixed-size window belong to the same segment) matches between the ground truth and the prediction. WD, on the other hand, measures whether the number of boundaries within the window is consistent with the ground truth. However, both metrics assume contiguity and are not suitable for evaluating non-contiguous text segmentation.
Therefore, we report Adjusted Rand Index (ARI) [11, 23] and Normalized Mutual Information (NMI) [13], which are appropriate for clustering-based evaluations. This adjustment accounts for the fact that the outputs of BP-Seg and $k$-means may result in non-contiguous segmentations, even though the ground truth segmentation is contiguous. Additionally, for efficiency, we use a variant of BP-Seg that is also based on message passing but offers faster computation. Please refer to Algorithm 2 for more details.
The ARI and Normalized NMI are two widely used metrics for evaluating clustering and segmentation quality. ARI ranges from 1-1 to $1$, where a score of $1$ indicates a perfect agreement between the predicted segmentation and the ground truth, 0 reflects a performance equivalent to random labeling, and negative values suggest an agreement worse than random chance. This makes ARI particularly informative in distinguishing between meaningful segmentations and those produced by chance. NMI, on the other hand, measures the mutual dependence between the predicted and true labels. It ranges from 0 to $1$, with $1$ indicating complete alignment between the predicted and true segmentations, and 0 representing statistical independence. Unlike ARI, NMI is insensitive to permutations of cluster labels, making it a complementary metric for evaluating clustering performance.
Table 1 reports the mean segmentation performance of the three methods on the Choi dataset. As illustrated in the table, BP-Seg consistently outperforms the baseline methods across nearly all configurations, except in cases where each segment contains very few sentences (e.g., 3–5). This highlights the strength of our belief propagation-based formulation in capturing both local and global textual coherence. The standard deviation of the performance can be read in Table 3.
@@ 5 Conclusion
We presented BP-Seg, an efficient unsupervised approach for text segmentation using belief propagation. Our method effectively balances local contextual coherence with global semantic similarity, enabling more meaningful and flexible segmentation of text. Although designed for non-contiguous segmentation, experimental results show that BP-Seg outperforms several competitive methods on the standard contiguous segmentation task, achieving strong performance on metrics such as ARI and NMI. Looking forward, we aim to explore the utility of BP-Seg in real-world downstream applications. These include prompt pruning for LLMs, where segmenting prompts into semantically coherent chunks can lead to more efficient inference; information retrieval, where flexible segmentation can support better indexing and matching; and question answering, where isolating relevant spans is crucial for interpretability and performance. We believe the probabilistic and modular nature of BP-Seg offers a promising foundation for broader integration into these complex language understanding pipelines.
@@
In this study, all examples are in English. The example in Section 4.1 was generated by GPT-4o with additional human-written content, and the Choi dataset in Section 4.2 is also synthetic. As a result, these examples may not accurately reflect real-world scenarios, and our evaluations are limited to these two cases. Nevertheless, we believe the insights from our findings will inspire further research in text segmentation and benefit a wide range of related applications.
@@ A.1.1 Results using BP-Seg
: [Segment 1]: The sun was shining brightly. It was a beautiful morning. I decided to go for a walk. Suddenly, dark clouds appeared. I’ll play tennis tomorrow. What are you doing? The rain poured down heavily. People ran for shelter. US Open is a tennis tournament. I am here working on my project. The sun came out again. Who is going to win the US Open? [Segment 2]: Thunder rumbled in the distance.
@@
$k=4$ segments: [Segment 1]: The sun was shining brightly. It was a beautiful morning. I decided to go for a walk. What are you doing? People ran for shelter. I am here working on my project. The sun came out again. [Segment 2]: Suddenly, dark clouds appeared. The rain poured down heavily. [Segment 3]: Thunder rumbled in the distance. [Segment 4]: I’ll play tennis tomorrow. US Open is a tennis tournament. Who is going to win the US Open?
$k=5$ segments: [Segment 1]: The sun was shining brightly. It was a beautiful morning. Suddenly, dark clouds appeared. Thunder rumbled in the distance. People ran for shelter. US Open is a tennis tournament. The sun came out again. [Segment 2]: I decided to go for a walk. [Segment 3]: I’ll play tennis tomorrow. What are you doing? Who is going to win the US Open? [Segment 4]: The rain poured down heavily. [Segment 5]: I am here working on my project.
$k=6$ segments: [Segment 1]: The sun was shining brightly. It was a beautiful morning. The rain poured down heavily. The sun came out again. [Segment 2]: I decided to go for a walk. [Segment 3]: I’ll play tennis tomorrow. US Open is a tennis tournament. Who is going to win the US Open? [Segment 4]: Suddenly, dark clouds appeared. Thunder rumbled in the distance. [Segment 5]: People ran for shelter. [Segment 6]: What are you doing? I am here working on my project.
@@ A.2 Choi dataset
Here $m_{i}(x_{i})$ can be viewed as a running preference or partial belief for node $i$ in label $x_{i}$. We also set $\psi_{i}(x_{i})=\mathrm{sim}(R_{i},C_{x_{i}})$ and $\psi_{i,j}=\lambda\mathrm{sim}(R_{i},R_{j})\exp(-|i-j|^{2}/\sigma)$. For comparison, we set the number of initial clusters in BP-Seg to be equal to the number of sentences, although many clusters may become empty after the algorithm terminates. We fix the number of iterations to $T=5$, and choose $\sigma=10$ and $\lambda=300$ to account for the length of the text. For GraphSeg, we set the threshold parameter to $\tau=0.2$ and the minimum segment size to $n=1$ to allow maximum flexibility. For the $k$-means baseline, we cap the number of clusters $k$ at 20 and rely on the default number of iterations in sklearn.cluster.KMeans.